OpenStudy (anonymous):

Not sure how to solve this. f(x) x^2+12x+27 g(x)=x+9 (f/g)(x)

OpenStudy (anonymous):

you have to solve this (x² + 12x +27)/(x+9), you can use many methods, briot ruffini and others

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i still don't know how to solve this

OpenStudy (anonymous):

\[(f/g)(x) = \frac{f(x)}{g(x)} = \frac{x^{2} + 12x + 27}{x + 9}\]

OpenStudy (anonymous):

Try factoring x^2+12x+27. If you get x+9 as one factor,then the other factor is the solution.

OpenStudy (anonymous):

i understand how to set it up but i ended up witht he wrong answer (x+2)

OpenStudy (anonymous):

Alright. First, factor \(x^{2} + 12x + 27\).Can you do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Find factors of 27 that add to 12. Then place the factors where I signify them: \[x^{2} + 12x + 27 = (x + \space factor1)(x + \space factor2)\]

OpenStudy (anonymous):

(x+3)(x+9)

OpenStudy (anonymous):

\[\small(f/g)(x) = \frac{f(x)}{g(x)} = \frac{x^{2} + 12x + 27}{x + 9} = \frac{(x + 9)(x + 3)}{x + 9} = \frac{\cancel{(x + 9)}(x + 3)}{\cancel{x + 9}} = x + 3\]

OpenStudy (anonymous):

i get it. thanks!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (anonymous):

Do you have it posted?

OpenStudy (anonymous):

f(x)=2x^2-4x+6 and g(x)=x-1; (f x g)(x)

OpenStudy (anonymous):

Just do: (x - 1)(2x^2 - 4x + 6) = ?

OpenStudy (anonymous):

2x^3-2x^2+10x+6

OpenStudy (anonymous):

Close. It's \(2x^{3} - 6x^{2} + 10x - 6\)

OpenStudy (anonymous):

those signs get me! could you help me figure out what i did on wrong on one more question?

OpenStudy (anonymous):

Alright. By your wording, I'm guessing you have an answer and you want me to check it?